OpenStudy (anonymous):

Find three solutions of the equation y = 9x – 4. (1 point) • (–5, –49), (–2, –22), (3, 23) • (–5, –49), (2, –22), (3, 23) • (–5, –49), (–2, 22), (–3, 23) • (5, –49), (–2, 22), (–3, 23)

OpenStudy (anonymous):

first horiz.points satisfy the equation -49=9(-5)-4 =-45- 4 =-49 just put in the x and y values as i have done

OpenStudy (anonymous):

would it be A